### v1

# Multi-Metric Preference Alignment for Generative Speech Restoration

Abstract
Recent generative models have significantly advanced speech restoration tasks, yet their training objectives often misalign with human perceptual preferences, resulting in suboptimal quality. While post-training alignment has proven effective in other generative domains like text and image generation, its application to generative speech restoration remains largely under-explored. This work investigates the challenges of applying preference-based post-training to this task, focusing on how to define a robust preference signal and curate high-quality data to avoid reward hacking. To address these challenges, we propose a multi-metric preference alignment strategy. We construct a new dataset, GenSR-Pref, comprising 80K preference pairs, where each chosen sample is unanimously favored by a complementary suite of metrics covering perceptual quality, signal fidelity, content consistency, and timbre preservation. This principled approach ensures a holistic preference signal. Applying Direct Preference Optimization (DPO) with our dataset, we observe consistent and significant performance gains across three diverse generative paradigms: autoregressive models (AR), masked generative models (MGM), and flow-matching models (FM) on various restoration benchmarks, in both objective and subjective evaluations. Ablation studies confirm the superiority of our multi-metric strategy over single-metric approaches in mitigating reward hacking. Furthermore, we demonstrate that our aligned models can serve as powerful “data annotators”, generating high-quality pseudo-labels to serve as a supervision signal for traditional discriminative models in data-scarce scenarios like singing voice restoration.
1 Introduction
Speech restoration, which aims to recover high-quality speech from various degradations, is a fundamental task in audio processing (Liu et al. 2022; Zhang et al. 2024b). Recent advances in utilizing generative models for speech restoration, namely generative speech restoration (GenSR), have shown remarkable performance in tasks such as denoising, dereverberation, declipping, super-resolution, etc. (Zhang et al. 2025a; Yang et al. 2024b; Li, Wang, and Liu 2024; Wang et al. 2025c, 2024b; Yao et al. 2025a; Kang et al. 2025; Wang et al. 2024a; Liu et al. 2024b). Unlike traditional discriminative methods that minimize a distance metric to a clean reference, generative speech restoration models are trained to maximize the likelihood of clean speech given a degraded input, thus learning the underlying distribution of clean speech (Lemercier et al. 2023), enabling them to generate high-fidelity audio even from severely degraded inputs.
In parallel, post-training techniques, designed to align generative models with specific downstream objectives or human preferences, have become integral to the advancement of generative modeling in some domains. In natural language processing (Ouyang et al. 2022; Rafailov et al. 2023; Bai et al. 2022), text-to-image/video (Xu et al. 2023; Wallace et al. 2024; Liu et al. 2025), and text-to-speech/audio (Zhang et al. 2025b; Sun et al. 2025; Hussain et al. 2025; Tian et al. 2025; Yao et al. 2025b; Zhang et al. 2024a; Liao et al. 2024) synthesis, preference-based alignment methods have been instrumental in enhancing model quality, safety, and human alignment. However, despite its potential, the application of such post-training alignment to the domain of generative speech restoration remains largely under-explored.
Applying preference-based post-training to GenSR models presents a unique set of challenges. Chief among them are: (1) Defining a faithful preference signal: How to construct an accessible, automated proxy that captures the multi-faceted nature of human auditory perception (which values clarity, naturalness, and a lack of artifacts)? (2) Curating high-quality preference data: Given a preference signal, what is an effective strategy for constructing preference pairs that can robustly guide model optimization? (3) Mitigating reward hacking: How to ensure the model achieves holistic, genuine improvement, rather than simply learning to exploit the biases of a specific metric?
To address these challenges, we propose and investigate a multi-metric preference alignment strategy. We argue that a robust solution to reward hacking lies in the preference signal itself being multi-dimensional and holistic. To this end, we construct a new preference dataset, which we name GenSR-Pref. A sample is chosen as winner in GenSR-Pref only when it is unanimously judged superior by a complementary suite of metrics that assess four distinct aspects of quality: perceptual quality, signal-level fidelity, content alignment, and timbre preservation. This strict, unanimous criterion ensures that our model learns from a robust and holistic preference signal, effectively mitigating the risk of reward hacking.
Based on the GenSR-Pref dataset, we apply Direct Preference Optimization (DPO) (Rafailov et al. 2023) to align models across three major distinct generative paradigms: the sequential Autoregressive (AR) models, the iterative Masked Generative (MGM) models (also known as masked or discrete diffusion), and the continuous Flow-Matching (FM) models. Our approach yields significant and consistent improvements across all benchmarks even with only 3k pairs, confirmed by objective and subjective tests. Our key findings are threefold. First, our multi-metric signal is crucial for avoiding the reward hacking inherent in single-metric optimization. Second, we show that using ground-truth as a fixed winner causes model collapse, highlighting the need for learned, relative preferences over absolute targets. Third, models perform best with data from their own architecture, a principle we term in-paradigm alignment. We attribute this to unique “alignment directions” for each paradigm, where in-paradigm data provides a more direct optimization path. Finally, we demonstrate that our aligned models can act as powerful “data annotators”, generating high-quality pseudo-labels to train discriminative models in data-scarce scenarios like singing voice restoration, thus bridging the gap between generative and discriminative paradigms.
To summarize, our main contributions are:
  • We propose a multi-metric preference alignment strategy to address the challenges of defining a robust preference signal and mitigating reward hacking in post-training for generative speech restoration. Through extensive experiments, we demonstrate that this strategy yields significant and consistent improvements across three diverse generative paradigms (AR, MGM, and FM).
  • We construct and introduce GenSR-Pref, a dataset of 80k preference pairs built upon a strict, unanimous multi-metric agreement criterion. This dataset, which captures a holistic aspect of audio quality, will be publicly released to facilitate future research.
  • We showcase a practical application of our aligned models, demonstrating their ability to serve as “data annotators” to generate pseudo-labels, thereby empowering the training of traditional discriminative models in data-scarce scenarios.
2 Related Work
2.1 Generative Speech Restoration
The paradigm for speech restoration has recently shifted towards generative modeling, which can be broadly categorized by the underlying generation mechanism. One major stream leverages autoregressive language models to treat restoration as a sequence-to-sequence problem, as seen in SELM (Wang et al. 2024b), TSELM (Tang, Zeng, and Li 2024), GenSE (Yao et al. 2025a), and versatile multi-task systems like SpeechX (Wang et al. 2024a), UniAudio (Yang et al. 2024a), and LLaSE-G1 (Kang et al. 2025). A second family consists of masked generative models, or discrete diffusion models, which iteratively refine tokenized audio representations. This approach has been applied to general restoration (e.g., MaskSR (Li, Wang, and Liu 2024; Liu et al. 2024c), AnyEnhance (Zhang et al. 2025a)), speech enhancement (e.g., Genhancer (Yang et al. 2024b)), and target speaker extraction (e.g., Metis (Wang et al. 2025c)). A third direction employs generative models with continuous dynamics to learn mappings in a continuous space; this includes methods based on score-SDE (Welker, Richter, and Gerkmann 2022), latent diffusion (Liu et al. 2024b; Wang et al. 2025a), flow-matching (Wang et al. 2025d; Liu et al. 2024a; Ku et al. 2024), and schrödinger bridges (Li et al. 2025a; Jukić et al. 2024) for various restoration tasks. Despite achieving state-of-the-art results, these generative models are typically trained using likelihood-based objectives, which often misalign with human perceptual quality. Our work addresses this gap by introducing a preference-based post-training strategy applicable across these diverse generative paradigms.
2.2 Post-Training for Audio Alignment
Post-training alignment has proven effective across various generative domains, including text (Ouyang et al. 2022; Rafailov et al. 2023; Bai et al. 2022), vision (Xu et al. 2023; Wallace et al. 2024; Liu et al. 2025), speech (Zhang et al. 2025b; Sun et al. 2025; Hussain et al. 2025; Tian et al. 2025; Yao et al. 2025b; Zhang et al. 2024a), music (Cideron et al. 2024; Lei et al. 2025), and audio effects (Majumder et al. 2024; Liao et al. 2024). In the context of speech restoration, early efforts like MetricGAN (Fu et al. 2019) used adversarial training to optimize for metrics like PESQ. More recently, preference-based methods have emerged, such as aligning with a learned NISQA predictor via PPO (Kumar, Perrault, and Williamson 2025) or using DPO with a single metric like UTMOS (Li et al. 2025b). While single-metric alignment is effective, it risks reward hacking by ignoring that robust restoration is multi-faceted, requiring a balance of quality, fidelity, and timbre (Section 4.3). Grounded in this principle, we propose a multi-metric alignment strategy to pursue holistic improvement. We systematically show this approach yields more comprehensive gains across diverse generative paradigms while naturally mitigating reward hacking.
3 Multi-Meric Preference Alignment
Figure 1: An overview of our multi-metric preference alignment strategy. The process consists of two main stages: (1) constructing the GenSR-Pref preference dataset by ranking model outputs based on a unanimous agreement across multiple metrics, and (2) fine-tuning the model with these preferences using Direct Preference Optimization (DPO).
Our approach to aligning generative speech restoration models is centered around a multi-metric preference alignment strategy, which is illustrated in Figure 1. This strategy involves two key stages: first, the construction of a high-quality preference dataset, and second, the application of Direct Preference Optimization (DPO) to align the model with these preferences. This approach is designed to be model-agnostic, allowing it to be applied to any generative restoration model.
3.1 The GenSR-Pref Dataset: Curating a Holistic Preference Signal
A pivotal element of our alignment strategy is the GenSR-Pref dataset, designed to provide a robust, holistic preference signal while mitigating single-metric reward hacking. To ensure our dataset’s breadth and quality, we generate candidate outputs using models from three distinct generative paradigms. We then form preference pairs by ranking outputs exclusively within each paradigm. This approach is chosen to capture the subtle variations in a single model’s generations, providing a fine-grained signal tailored to its specific error patterns. This comprehensive strategy yields a dataset for robust alignment and enables detailed in-paradigm versus cross-paradigm analyses.
Model Selection
To comprehensively evaluate our alignment framework across different generative paradigms, we adopt one representative model from each of the three major families: autoregressive (AR), masked generative (MGM), and flow-matching (FM) models.
Masked Generative Model (MGM). We use the pre-trained AnyEnhance (Zhang et al. 2025a) model as our MGM model. It predicts acoustic tokens based on the DAC codec (Kumar et al. 2024) from partially masked sequences, effectively capturing non-autoregressive and context-aware dependencies.
Autoregressive Model (AR). We train a new model, AR+Soundstorm, which follows a two-stage pipeline common in text-to-speech (TTS) domain. It first predicts semantic tokens of the Metis codec (Wang et al. 2025c) given noisy inputs, and a second-stage non-autoregressive model similar to Soundstorm (Borsos et al. 2023) then converts these tokens into acoustic tokens for waveform synthesis.
Flow-Matching Model (FM). We train Flow-SR, a model that learns to predict the velocity field of clean mel-spectrograms from noisy inputs via optimal transport interpolation based on the Diffusion Transformer (DiT) architecture (Peebles and Xie 2023). A pre-trained Vocos vocoder (Siuzdak 2023) is then used to synthesize the final waveform from the generated spectrograms. Details on model architectures and training procedures can be found in Appendix B.
To ensure a reliable preference signal that captures the multi-faceted nature of audio quality, we evaluate each candidate output across four complementary dimensions:
  • Perceptual Quality (NISQA) (Mittag et al. 2021): Assesses overall listening quality, including naturalness and the absence of annoying artifacts.
  • Signal-level Fidelity (DNSMOS) (Reddy, Gopal, and Cutler 2022): A composite metric evaluating signal distortion, background noise, and overall quality for a fine-grained assessment of signal integrity.
  • Content Alignment (SpeechBERTScore) (Saeki et al. 2024): Measures semantic similarity to the ground-truth transcription to ensure linguistic content is not altered during restoration.
  • Timbre Preservation (Speaker Similarity): Uses a pre-trained speaker verification model to compute cosine similarity, ensuring the speaker’s identity is preserved.
A preference pair $(y_{w},y_{l})$ is formed only under a strict, unanimous agreement criterion: the winning output $y_{w}$ must score higher than the losing output $y_{l}$ on all of these metrics simultaneously. This process ensures each preference pair represents a holistic and unambiguous improvement, providing a robust signal for the DPO alignment stage.
In total, GenSR-Pref comprises approximately 80K preference pairs, structured for distinct experimental goals. The primary component is a large-scale subset of 69,456 pairs from the AnyEnhance (MGM) model, allowing us to validate our alignment strategy’s effectiveness with a substantial amount of data. For fair comparisons and ablation studies—particularly our cross-paradigm analysis—we created smaller, controlled subsets from a shared set of input utterances, yielding 3,354 pairs for AR, 3,428 for FM, and 3,035 for MGM. This dual-component structure enables both an investigation into data scaling effects and rigorous, controlled analyses of alignment dynamics.
3.2 GenSR Model Alignment via Direct Preference Optimization
In the second stage of our approach, we align the generative speech restoration models using the GenSR-Pref dataset. We employ Direct Preference Optimization (DPO) (Rafailov et al. 2023), a simple yet powerful technique that directly optimizes a policy to satisfy preferences without explicit reward modeling or reinforcement learning. The core idea of DPO is to re-parameterize the reward function in the standard RLHF objective using the optimal policy, which leads to a simple contrastive loss over preferred and dispreferred samples. In this section, we detail how we adapt and apply DPO to the three diverse generative paradigms used in our study: Autoregressive (AR), Masked Generative (MGM), and Flow-Matching (FM) models. Detailed derivations can be found in Appendix A.
DPO for Autoregressive Models
AR models treat speech restoration as a sequence-to-sequence task. Given a degraded input $x$, they generate a clean audio sequence $y=(y_{1},...,y_{T})$ token by token. They are typically trained by maximizing the log-likelihood of the ground-truth sequence, using a standard cross-entropy loss:
$\mathcal{L}_{\text{AR}}=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}\left[\sum_{t=1}^{T}\log p_{\theta}(y_{t}|y_{<t},x)\right]$ (1)
DPO Derivation. The standard RLHF objective aims to find a policy $p_{\theta}$ that maximizes the reward from a learned reward model $r_{\phi}(x,y)$, while constraining its deviation from a reference policy $p_{\text{ref}}$ using a KL-divergence term:
$\max_{p_{\theta}}\,\mathbb{E}_{y\sim p_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta D_{\text{KL}}\left[p_{\theta}(y|x)\,\|\,p_{\text{ref}}(y|x)\right]$ (2)
The key insight of DPO is that for this objective, the optimal policy $p^{*}_{\theta}$ has a closed-form solution: $p^{*}_{\theta}(y|x)\propto p_{\text{ref}}(y|x)\exp(\frac{1}{\beta}r_{\phi}(x,y))$. By rearranging this equation, the unknown reward function $r_{\phi}(x,y)$ can be expressed in terms of the optimal policy and the reference policy. Substituting this back into the Bradley-Terry loss for the reward model, which models the probability of preferring $y_{w}$ over $y_{l}$, yields the final DPO loss that directly optimizes the policy $p_{\theta}$:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigg{[}\log\sigma\Big{(}\beta\log\frac{p_{\theta}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}-\beta\log\frac{p_{\theta}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\Big{)}\Bigg{]}$ (3)
where $\mathcal{D}$ is our GenSR-Pref preference dataset, and $\sigma$ is the sigmoid function.
DPO for Masked Generative Models
MGM models (Chang et al. 2022; Zhang et al. 2025a; Wang et al. 2025c; Ju et al. 2024; Wang et al. 2025b), also known as discrete diffusion models, learn to restore a clean speech sequence $y_{0}$ from a partially masked or corrupted version $y_{t}$, where $t\in[0,T]$ is the noise step parameter that controls the amount of corruption, and $y_{T}$ is the fully masked sequence. They are trained to predict the original tokens in the masked positions, optimizing a masked language modeling objective:
$\mathcal{L}_{\text{MGM}}=-\mathbb{E}_{(x,y_{0}),t,m_{t}}\left[\sum_{i=1}^{T}m_{t,i}\log p_{\theta}(y_{0,i}|y_{t},x)\right]$ (4)
where $m_{t}\in\{0,1\}^{T}$ is a binary mask indicating which tokens in the sequence are masked (1) or unmasked (0). The model learns to predict the original tokens in the masked positions, effectively learning to denoise the corrupted input.
DPO Derivation. Following the INTP framework (Zhang et al. 2025b), we extend DPO to this non-autoregressive setting. The core logic remains the same, but the policy now represents the conditional probability of the full clean sequence $y_{0}$ given the corrupted version $y_{t}$. The DPO loss for MGM is thus analogous to the AR case, contrasting the log-probabilities of the preferred and dispreferred sequences, conditioned on their respective masked inputs:
$\mathcal{L}_{\text{DPO-MGM}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D},\,t}\Bigg{[}\log\sigma\Big{(}\beta\log\frac{p_{\theta}(y_{0}^{w}|y_{t}^{w},x)}{p_{\text{ref}}(y_{0}^{w}|y_{t}^{w},x)}$ (5)
${}-\beta\log\frac{p_{\theta}(y_{0}^{l}|y_{t}^{l},x)}{p_{\text{ref}}(y_{0}^{l}|y_{t}^{l},x)}\Big{)}\Bigg{]}$
Here, $y_{t}^{w}$ and $y_{t}^{l}$ are the masked versions of the winning ($y_{0}^{w}$) and losing ($y_{0}^{l}$) sequences from our preference dataset.
DPO for Flow-Matching Models
FM models (Lipman et al. 2022) learn a continuous mapping from a simple prior distribution (e.g., Gaussian noise $y_{0}\sim\mathcal{N}(0,I)$) to a complex data distribution (e.g., clean speech $y_{1}$). They are trained to predict the velocity vector field $v_{\theta}(y_{t},t,x)$ that describes the probabilistic flow of data points along a linear interpolation path from the noise $y_{0}$ to the data $y_{1}$. The standard training objective minimizes the difference between the predicted velocity and the true velocity:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t,y_{0},y_{1},x}\left[\|v_{\theta}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}\right]$ (6)
where $y_{t}=(1-t)y_{0}+ty_{1}$ is a point on the linear trajectory from the noise to the data, and $t$ is a noise step parameter that controls the interpolation between the two. The model learns to predict the velocity field that describes the continuous flow of data points along the linear interpolation path, effectively learning to denoise by predicting the gradient of the data distribution along the linear interpolation.
DPO Derivation. To extend DPO to FM models, we follow recent advances in aligning diffusion-based generators (Wallace et al. 2024; Liu et al. 2025). Instead of modeling the full data likelihood along the entire generative trajectory, we use a tractable surrogate objective that operates on a single noise step. Specifically, we approximate the DPO objective using the model’s instantaneous prediction error at a sampled timestep, treating the L2 loss between predicted and true velocities as a proxy for negative log-likelihood. The resulting loss encourages the model to reduce velocity error for preferred samples ($y_{w}$) while increasing it for dispreferred ones ($y_{l}$), yielding the following expression:
$\mathcal{L}_{\text{DPO-FM}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D},t,y_{0}}\left[\log\sigma\left(-\beta(\Delta_{w}-\Delta_{l})\right)\right]$ (7)
where $\Delta$ represents the difference in squared L2 error between the trained model $v_{\theta}$ and the reference model $v_{\text{ref}}$ for a given sample ($y_{w}$ or $y_{l}$):
$\Delta_{w}=\|v_{\theta}(y_{t}^{w},t,x)-(y_{1}^{w}-y_{0})\|_{2}^{2}-\|v_{\text{ref}}(y_{t}^{w},t,x)-(y_{1}^{w}-y_{0})\|_{2}^{2}$ (8)
and $\Delta_{l}$ is defined analogously for the losing sample $y_{l}$. This single-step formulation provides a scalable and effective approximation to DPO in diffusion processes.
4 Experiments
4.1 Experimental Setup
Models and Datasets. We evaluate our multi-metric preference alignment strategy across the three generative paradigms: AnyEnhance for MGM, AR+Soundstorm for AR, and Flow-SR for FM. All alignment experiments are conducted using our newly constructed GenSR-Pref dataset. Specifically, we conduct the main experiment (Section 4.2) using the primary preference subset (69k pairs for MGM, approx. 3k for AR/FM), while the ablations (Section 4.3) are conducted using the controlled smaller 3k-pair subsets for all models to ensure fair and direct comparisons.
Dataset Model Type
DPO-
aligned?
[…]
Speech-
[…]
Voicefixer-GSR Voicefixer DISC. - 3.299 3.971 3.003 4.160 0.797 0.882
MaskSR MGM - 3.445 3.971 3.128 - - -
AnyEnhance MGM 3.406 4.073 3.136 4.308 0.829 0.924
3.532 4.091 3.267 4.639 0.834 0.935
AR+Soundstorm AR 3.550 4.097 3.294 4.556 0.788 0.894
3.564 4.144 3.331 4.850 0.803 0.904
Flow-SR FM 3.398 3.969 3.104 4.010 0.812 0.918
3.483 4.092 3.230 4.672 0.830 0.924
Librivox-GSR AnyEnhance MGM 3.546 4.142 3.308 4.346 0.822 0.955
3.690 4.201 3.475 4.865 0.828 0.954
AR+Soundstorm AR 3.663 4.134 3.419 4.535 0.783 0.922
3.693 4.193 3.478 4.923 0.793 0.924
Flow-SR FM 3.550 4.062 3.281 4.184 0.791 0.931
3.602 4.152 3.368 4.825 0.801 0.930
CCMusic-GSR AnyEnhance MGM 3.243 3.547 2.797 3.345 0.811 0.915
3.440 3.827 3.062 4.154 0.817 0.909
AR+Soundstorm AR 3.378 3.693 2.956 3.948 0.710 0.854
3.460 3.865 3.094 4.438 0.713 0.853
Flow-SR FM 3.298 3.540 2.813 3.897 0.733 0.883
3.378 3.770 2.971 4.371 0.740 0.881
Table 1: Quantitative comparison on GSR benchmarks across three generative paradigms—MGM (AnyEnhance), AR (AR+Soundstorm), and FM (Flow-SR)—before and after DPO alignment. Results from a discriminative baseline Voicefixer (DISC.) and the MaskSR model are also included. Bold indicates improvements after alignment.
Evaluation Benchmarks and Metrics. We conduct evaluations on two major kinds of benchmarks: (1) General speech restoration (GSR) that includes denoising, dereveration, declipping and super-resolution at the same time (Voicefixer-GSR, Librivox-GSR, CCMusic-GSR), (2) Speech Enhancement (SE) under noisy and reverberant conditions (DNS-No-Reverb and DNS-With-Reverb). For a comprehensive evaluation, we not only include the original unaligned models, but also involve state-of-the-art baselines LLaSE-G1 (Kang et al. 2025), GenSE (Yao et al. 2025a), MaskSR (Li, Wang, and Liu 2024), and FlowSE (Wang et al. 2025d), and a discriminative model Voicefixer (Liu et al. 2022). Performance is measured using a comprehensive suite of objective metrics: DNSMOS (SIG, BAK, OVRL) (Reddy, Gopal, and Cutler 2022) for signal fidelity and quality, NISQA (Mittag et al. 2021) for perceptual quality, SpeechBERTScore (SBERT) (Saeki et al. 2024) for content consistency, and speaker similarity (SIM). For subjective evaluation, we conduct A/B preference tests to assess human perception. Details on the evaluation can be found in Appendix C.
4.2 Main Results: Effectiveness of Multi-Metric Preference Alignment
Table 1 presents the main results, showing that our multi-metric preference alignment strategy is consistently effective across all three paradigms. Notably, the MGM model achieves significant gains when aligned on its large-scale 69k-pair subset (e.g., +0.519 NISQA on LibriVox-GSR), demonstrating our method’s scalability. Impressively, the AR and FM models also show significant improvements (e.g., +0.388 and +0.641 NISQA) even with their much smaller, 3k-pair data, highlighting the data efficiency of our approach. These objective gains are confirmed by subjective A/B tests (Figure 2), where human listeners consistently preferred the aligned models, with win rates reaching up to 54.5%.
Figure 2: Human AB preference test results for DPO-aligned models on Librivox-GSR testset.
Moreover, this strong performance generalizes beyond GSR tasks. As shown in Table 2, models aligned with our diverse preference data also exhibit seamless performance gains on the downstream DNS speech enhancement benchmark. For example, the aligned AnyEnhance model boosts its OVRL score from 3.204 to 3.438 on the reverberant set, demonstrating the robustness of our alignment strategy. More results and audio visualizations can be found in Appendix D.
Data Model Type Aligned? SIG BAK OVRL
No
Reverb
MaskSR MGM - 3.586 4.116 3.339
GenSE AR - 3.650 4.180 3.430
LLaSE-G1 AR - 3.660 4.170 3.420
FlowSE FM - 3.685 4.201 3.445
AnyEnhance MGM 3.640 4.179 3.418
3.684 4.203 3.476
AR+Soundstorm AR 3.648 4.155 3.422
3.673 4.194 3.469
Flow-SR FM 3.581 4.133 3.355
3.632 4.173 3.420
With
Reverb
MaskSR MGM - 3.531 4.065 3.253
GenSE AR - 3.490 3.730 3.190
LLaSE-G1 AR - 3.590 4.100 3.330
FlowSE FM - 3.601 4.102 3.331
AnyEnhance MGM 3.500 4.040 3.204
3.670 4.178 3.438
AR+Soundstorm AR 3.681 4.127 3.431
3.709 4.189 3.496
Flow-SR FM 3.539 4.019 3.255
3.629 4.163 3.399
Table 2: Evaluation results on the DNS speech enhancement benchmark. Models aligned on diverse restoration preference pairs (e.g., denoising, declipping) exhibit seamless performance gains when applied to speech enhancement as a downstream task.
4.3 Ablation Studies and Analysis
To better understand the factors contributing to our performance gains, we conduct a series of in-depth ablation studies.
Multi-Metric vs. Single-Metric Alignment. To validate our core hypothesis that a multi-metric signal is crucial for avoiding reward hacking, we compare our Multi-Metric Preference Alignment approach with DPO alignment using preference pairs constructed from single metrics. As shown in Table 3, while optimizing for a single metric improves that specific score, it often leads to stagnation or even degradation in other, un-targeted metrics. For example, the Similarity-aligned model performs worse than the baseline on OVRL and SIG. In contrast, our multi-metric approach achieves robust improvements across all metrics simultaneously. This confirms that our unanimous agreement criterion is effective at mitigating reward hacking and promoting holistic quality improvement. More results can be found in Appendix D.3.
Criterion SIG BAK OVRL NISQA SBERT SIM
- 3.550 4.097 3.294 4.556 0.788 0.894
Multi-Metric 3.564 4.144 3.331 4.850 0.803 0.904
NISQA 3.531 4.137 3.300 4.810 0.785 0.896
OVRL 3.561 4.117 3.317 4.600 0.792 0.896
SIM 3.537 4.101 3.285 4.577 0.792 0.901
SBERT 3.540 4.109 3.291 4.612 0.804 0.901
Table 3: Ablation on preference metric for the AR model on Voicefixer-GSR testset. The Multi-Metric Preference Strategy (”Multi-Metric”) consistently outperforms single-metric approaches across all evaluation metrics.
(a) SIG
(b) BAK
(c) OVRL
(d) NISQA
(e) SpeechBERTScore
(f) Similarity
Figure 3: Ablation study on training objectives for the MGM model. DPO demonstrates consistent improvements across training steps, while SFT tends to stagnate or degrade. Notably, a naive DPO variant that treats ground-truth outputs as winners (”GT Winner”) results in model collapse.
(a) Reward Margin ($\times 10^{5}$)
(b) Reward Accuracy
Figure 4: Detailed training curves between normal DPO and DPO (GT Winner) training (smoothing factor = 0.99). Using ground truth as the unconditional winner leads to inflated reward margins and saturated reward accuracy, indicating model collapse.
DPO vs. Supervised Fine-Tuning (SFT). To highlight the benefits of preference-based optimization, we compare DPO against two SFT baselines: fine-tuning on ground-truth audio (SFT-GT) and on the ”winner” samples from our preference data (SFT-Winner). As shown in Figure 3, DPO consistently outperforms both SFT variants, which show only marginal gains before stagnating. This suggests that simply exposing the model to high-quality examples is insufficient for effective alignment. Furthermore, a naive DPO strategy using ground-truth as the fixed “winner” leads to model collapse, as detailed in Figure 4. This behavior is driven by the core DPO objective, which seeks to maximize the log-probability ratio, $\beta\log\frac{p_{\theta}(y_{0}^{w}|y_{t}^{w},x)}{p_{\text{ref}}(y_{0}^{w}|y_{t}^{w},x)}-\beta\log\frac{p_{\theta}(y_{0}^{l}|y_{t}^{l},x)}{p_{\text{ref}}(y_{0}^{l}|y_{t}^{l},x)}$. When $y_{w}$ is a fixed ground-truth target, the model learns a pathological shortcut—drastically suppressing the probability of all outputs other than the GT to maximize this ratio. This highlights the necessity of learning from relative, nuanced preferences.
Model (Source) SIG BAK OVRL NISQA SBERT SIM
AR (-) 3.663 4.134 3.419 4.535 0.783 0.922
AR (AR) 3.672 4.171 3.449 4.867 0.791 0.925
AR (MGM) 3.675 4.148 3.435 4.604 0.786 0.919
AR (FM) 3.668 4.166 3.440 4.783 0.788 0.918
FM (-) 3.550 4.062 3.281 4.184 0.791 0.931
FM (AR) 3.575 4.103 3.324 4.473 0.794 0.932
FM (MGM) 3.587 4.048 3.305 4.269 0.789 0.930
FM (FM) 3.580 4.126 3.337 4.645 0.797 0.929
MGM (-) 3.546 4.142 3.308 4.346 0.822 0.955
MGM (AR) 3.589 4.147 3.346 4.321 0.809 0.951
MGM (MGM) 3.644 4.155 3.409 4.652 0.825 0.958
MGM (FM) 3.478 4.154 3.232 4.468 0.802 0.946
Table 4: Ablation on the source of preference data, evaluated on the Librivox-GSR testset. For each base model, performance is compared against its baseline (-) when aligned with data from different model paradigms (AR, MGM, FM).
AR FM MGM
AR 0.0075 0.0072 0.0018
FM 0.0072 0.0102 0.0005
MGM 0.0018 0.0005 0.0045
Table 5: Average pairwise cosine similarity of preference vectors ($v_{\text{pref}}=f(y_{w})-f(y_{l})$) within and across generative paradigms. Higher values indicate greater alignment in the ‘alignment direction’. The strong diagonal highlights that in-paradigm preference data is more internally consistent.
The Role of Data Source in Alignment. Finally, we explore how the source of preference data affects alignment performance. Table 4 shows the results of aligning each model with preference pairs generated from itself (in-paradigm) versus from other paradigms (cross-paradigm). While models can often benefit from cross-paradigm data, we observe a clear trend: all three models achieve their optimal performance when aligned with their own in-paradigm preference data. To quantitatively investigate this phenomenon, we formally define the principle of in-paradigm alignment. We hypothesize that each architecture possesses a unique “preference direction”—an optimal path for improvement. To quantify this, we define a “preference vector” as the feature-space difference between winner and loser samples ($v_{\text{pref}}=f(y_{w})-f(y_{l})$) using pre-trained w2v-bert-2 model [^2] https://huggingface.co/facebook/w2v-bert-2.0 and compute their average pairwise cosine similarity. The results (Table 5) provide strong quantitative evidence for our hypothesis: the average in-paradigm similarity (diagonal values) is consistently higher than the cross-paradigm similarity (e.g., AR-AR similarity of 0.0075 vs. AR-MGM of 0.0018). This indicates that data from a model’s own paradigm provides a more consistent and targeted optimization signal.
Furthermore, these similarities correlate with alignment performance (Table 4): paradigms with more aligned preference vectors (e.g., AR and FM, similarity 0.0072) show better cross-paradigm transfer performance than less aligned ones (e.g., FM and MGM, similarity 0.0005). This analysis provides strong backing for our principle, suggesting that alignment is most effective when the preference data’s direction matches the target model’s intrinsic generative manifold.
4.4 Application: Empowering Discriminative Models via Pseudo-Labeling
To showcase the practical utility of our aligned models, we explore their potential as “data annotators” in a scenario where ground-truth is scarce: real-world singing voice restoration. We took a standard discriminative restoration model, Voicefixer, and fine-tuned it on a set of singing recordings. The key difference is that the “clean” supervision signals were not ground-truth recordings, but pseudo-labels generated by our DPO-enhanced AnyEnhance model.
Model SIG BAK OVRL NISQA
Voicefixer (before) 2.657 3.080 2.295 2.919
Voicefixer (after) 3.096 3.745 2.756 3.312
Table 6: Evaluation on real-world singing-voice recordings before and after DPO model’s annotation
As shown in Table 6, the fine-tuned Voicefixer model achieves dramatic improvements across all metrics compared to its original state. This result is significant: it demonstrates that a powerful, DPO-aligned generative model can effectively create high-quality supervision signals, enabling smaller, more efficient discriminative models to be trained for tasks where obtaining paired data is prohibitively difficult. This bridges the gap between the two modeling paradigms and opens up new avenues for practical speech restoration. More detail on this application can be found in the Appendix D.4.
5 Conclusion
In this work, we introduced a multi-metric preference alignment strategy to align generative speech restoration models with human perception while mitigating the risk of reward hacking. To this end, we constructed a new dataset, GenSR-Pref, using a strict unanimous agreement criterion across four complementary metrics to ensure a holistic and robust preference signal. Applying DPO with our dataset yields significant improvements across diverse generative paradigms (AR, MGM, and FM), outperforming single-metric preference and demonstrating the principle of in-paradigm alignment. Furthermore, we demonstrated a novel and practical application where our aligned models serve as powerful “data annotators”, generating high-quality pseudo-labels to effectively train discriminative models in data-scarce scenarios. Future work includes exploring more advanced alignment algorithms, investigating preference data properties, and leveraging our annotation capabilities for large-scale data curation in other domains like TTS.
Appendix A Derivation of DPO Loss for Three Generative Paradigms
This section provides a detailed derivation of the Direct Preference Optimization (DPO) loss functions for the three generative paradigms explored in our work: Autoregressive (AR), Masked Generative Models (MGM), and Flow-Matching (FM). We begin with a summary table (Table 7) that juxtaposes the key components of DPO across these paradigms, followed by step-by-step derivations for each.
Paradigm RL Objective (with KL Regularization) Implicit Reward Mapping Final DPO Loss Formulation
AR $\max_{p_{\theta}}\,\mathbb{E}_{y\sim p_{\theta}(y|x)}[r(x,y)]-\beta D_{\text{KL}}\left[p_{\theta}(y|x)\,\|\,p_{\text{ref}}(y|x)\right]$ $\begin{aligned} r(x,y)&=\beta\log\frac{p^{*}_{\theta}(y|x)}{p_{\text{ref}}(y|x)}\\ &\quad+\beta\log Z(x)\end{aligned}$ $\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{p_{\theta}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}-\beta\log\frac{p_{\theta}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\right)\right]$
MGM $\max_{p_{\theta}}\,\mathbb{E}_{y_{0},t,x}[r(y_{0},x)]-\beta D_{\text{KL}}\left[p_{\theta}(y_{0}|y_{t},x)\,\|\,p_{\text{ref}}(y_{0}|y_{t},x)\right]$ $\begin{aligned} r(y_{0},x)&=\beta\log\frac{p^{*}_{\theta}(y_{0}|y_{t},x)}{p_{\text{ref}}(y_{0}|y_{t},x)}\\ &\quad+\beta\log Z(y_{t},x)\end{aligned}$ $\begin{aligned} \mathcal{L}_{\text{DPO-MGM}}=-\mathbb{E}_{\mathcal{D},t}\left[\log\sigma\left(\beta\log\frac{p_{\theta}(y_{0}^{w}|y_{t}^{w},x)}{p_{\text{ref}}(y_{0}^{w}|y_{t}^{w},x)}\right.\right.\\ \left.\left.\quad-\beta\log\frac{p_{\theta}(y_{0}^{l}|y_{t}^{l},x)}{p_{\text{ref}}(y_{0}^{l}|y_{t}^{l},x)}\right)\right]\end{aligned}$
FM $\max_{p_{\theta}}\,\mathbb{E}_{y_{1},t,x}[r(y_{1},x)]-\beta D_{\text{KL}}\left[p_{\theta}(y_{1}|y_{t},t,x)\,\|\,p_{\text{ref}}(y_{1}|y_{t},t,x)\right]$ $\begin{aligned} r(y_{1},x)&=\beta\log\frac{p^{*}_{\theta}(y_{1}|y_{t},t,x)}{p_{\text{ref}}(y_{1}|y_{t},t,x)}\\ &\quad+\beta\log Z(y_{t},t,x)\end{aligned}$ $\begin{aligned} \mathcal{L}_{\text{DPO-FM}}=-\mathbb{E}_{\mathcal{D},t}\log\sigma\left(-\beta(\Delta_{w}-\Delta_{l})\right),\\ \text{where}\quad\Delta=\|v_{\theta}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}\\ \quad-\|v_{\text{ref}}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}\end{aligned}$
Table 7: Direct Preference Optimization (DPO) across different generative paradigms. The DPO framework is adapted from standard Autoregressive (AR) models to Masked Generative Models (MGM) and Flow-Matching (FM) models. This table summarizes the RL objective, implicit reward mapping, and final DPO loss formulation.
A.1 DPO for Autoregressive Models
The derivation for AR models follows the original DPO framework (Rafailov et al. 2023). The goal is to optimize a policy $p_{\theta}(y|x)$ to maximize a reward function $r(x,y)$, while regularizing its deviation from a reference policy $p_{\text{ref}}(y|x)$ using a KL-divergence term. The RL objective is:
$\max_{p_{\theta}}\,\mathbb{E}_{y\sim p_{\theta}(y|x)}[r(x,y)]-\beta D_{\text{KL}}\left[p_{\theta}(y|x)\,\|\,p_{\text{ref}}(y|x)\right].$ (9)
The optimal policy that maximizes the RL objective is:
$p^{*}_{\theta}(y|x)=\frac{1}{Z(x)}p_{\text{ref}}(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right),$ (10)
where $Z(x)$ ensures normalization. Thus, the reward can be written as:
$r(x,y)=\beta\log\frac{p^{*}_{\theta}(y|x)}{p_{\text{ref}}(y|x)}+\beta\log Z(x).$ (11)
DPO re-parameterizes the reward model using this relationship and substitutes it into the binary cross-entropy loss for a preference dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$, where $y_{w}$ is preferred over $y_{l}$. This results in a simple maximum likelihood objective that directly optimizes the policy $p_{\theta}$:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigg{[}\log\sigma\Big{(}\beta\log\frac{p_{\theta}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}{}$ (12)
$-\beta\log\frac{p_{\theta}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\Big{)}\Bigg{]}$
A.2 DPO for Masked Generative Models (MGM)
The RL objective is similar to that of AR models, but the policy to optimize is conditioned on a masked input $y_{t}$ and a context $x$:
$\max_{p_{\theta}}$ $\mathbb{E}_{y_{0}\sim p_{\theta}(y_{0}|x),\,t,\,x}\left[r(y_{0},x)\right]$ (13)
${}-\beta\,\text{D}_{\text{KL}}\left[p_{\theta}(y_{0}|y_{t},x)\,\|\,p_{\text{ref}}(y_{0}|y_{t},x)\right].$
The optimal policy takes the form:
$p^{*}_{\theta}(y_{0}|y_{t},x)=\frac{1}{Z(y_{t},x)}p_{\text{ref}}(y_{0}|y_{t},x)\exp\left(\frac{1}{\beta}r(y_{0},x)\right),$ (14)
and the reward becomes:
$r(y_{0},x)=\beta\log\frac{p^{*}_{\theta}(y_{0}|y_{t},x)}{p_{\text{ref}}(y_{0}|y_{t},x)}+\beta\log Z(y_{t},x).$ (15)
Substituting yields the DPO loss for MGM:
$\mathcal{L}_{\text{DPO-MGM}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D},\,t}\Bigg{[}\log\sigma\Big{(}\beta\log\frac{p_{\theta}(y_{0}^{w}|y_{t}^{w},x)}{p_{\text{ref}}(y_{0}^{w}|y_{t}^{w},x)}$ (16)
${}-\beta\log\frac{p_{\theta}(y_{0}^{l}|y_{t}^{l},x)}{p_{\text{ref}}(y_{0}^{l}|y_{t}^{l},x)}\Big{)}\Bigg{]}$
A.3 DPO for Flow-Matching Models
For Flow-Matching (FM) models, the RL objective is structured similarly while incorporating the flow-matching context $y_{t}$ and time step $t$ and the input $x$:
$\max_{p_{\theta}}$ $\mathbb{E}_{y_{1}\sim p_{\theta}(y_{1}|x),t,x}\left[r(y_{1},x)\right]$ (17)
${}-\beta\,\text{D}_{\text{KL}}\left[p_{\theta}(y_{1}|y_{t},t,x)\,\|\,p_{\text{ref}}(y_{1}|y_{t},t,x)\right].$
The optimal policy:
$p^{*}_{\theta}(y_{1}|y_{t},t,x)=\frac{1}{Z(y_{t},t,x)}p_{\text{ref}}(y_{1}|y_{t},t,x)\exp\left(\frac{1}{\beta}r(y_{1},x)\right),$ (18)
with reward:
$r(y_{1},x)=\beta\log\frac{p^{*}_{\theta}(y_{1}|y_{t},t,x)}{p_{\text{ref}}(y_{1}|y_{t},t,x)}+\beta\log Z(y_{t},t,x).$ (19)
Because FM training assumes a Gaussian likelihood, following practice in (Wallace et al. 2024; Liu et al. 2025), we can express the policies into velocity predictions $v_{\theta}$ and $v_{\text{ref}}$ as follows:
$p_{\theta}(y_{1}|y_{t},t,x)$ $\propto\exp\left(-\frac{1}{\beta}\|v_{\theta}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}\right),$ (20)
$p_{\text{ref}}(y_{1}|y_{t},t,x)$ $\propto\exp\left(-\frac{1}{\beta}\|v_{\text{ref}}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}\right).$
Then we have the log-ratio:
$\log\frac{p_{\theta}(y_{1}|y_{t},t,x)}{p_{\text{ref}}(y_{1}|y_{t},t,x)}=-\frac{1}{\beta}\Big{(}$ $\|v_{\theta}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}$ (21)
$-\|v_{\text{ref}}(y_{t},t,x)-(y_{1}-y_{0})\|_{2}^{2}\Big{)}.$
Plugging into the DPO loss gives:
$\mathcal{L}_{\text{DPO-FM}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D},t,y_{0}}\left[\log\sigma\left(-\beta(\Delta_{w}-\Delta_{l})\right)\right],$ (22)
where
$\Delta_{w}=\|v_{\theta}(y_{t}^{w},t,x)-(y_{1}^{w}-y_{0})\|_{2}^{2}-\|v_{\text{ref}}(y_{t}^{w},t,x)-(y_{1}^{w}-y_{0})\|_{2}^{2}$ (23)
and $\Delta_{l}$ is defined analogously for the losing sample $y_{l}$. This loss encourages the model to reduce its velocity prediction error for preferred samples (making $\Delta_{w}$ more negative) while allowing for a larger error on dispreferred samples (making $\Delta_{l}$ less negative or positive) relative to the reference model.
Dataset Model Type
DPO-
aligned?
[…]
Speech-
[…]
DNS-No-Reverb AnyEnhance MGM 3.640 4.179 3.418 4.821 0.907 0.988
3.684 4.203 3.476 5.011 0.904 0.987
AR+Soundstorm AR 3.648 4.155 3.422 4.743 0.866 0.971
3.673 4.194 3.469 4.974 0.868 0.970
Flow-SR FM 3.581 4.133 3.355 4.711 0.893 0.984
3.632 4.173 3.420 4.998 0.895 0.983
DNS-With-Reverb AnyEnhance MGM 3.500 4.040 3.204 3.722 0.738 0.951
3.670 4.178 3.438 4.624 0.733 0.957
AR+Soundstorm AR 3.681 4.127 3.431 4.572 0.726 0.943
3.709 4.189 3.496 4.957 0.715 0.940
Flow-SR FM 3.539 4.019 3.255 4.215 0.744 0.955
3.629 4.163 3.399 4.924 0.746 0.955
Table 8: Quantitative comparison on DNS benchmarks across three generative paradigms before and after DPO alignment.
Appendix B Implementation Details
B.1 Model Details
Masked Generative Model (MGM). We use AnyEnhance as our model for MGM paradigm. AnyEnhance (Zhang et al. 2025a) is a unified masked generative model for speech and singing voice restoration, capable of handling a wide range of tasks (e.g., denoising, dereverberation, super-resolution) via prompt-guided in-context learning and a self-critic sampling mechanism. It is based on a transformer architecture (Vaswani et al. 2017) trained with a masked language modeling objective, allowing the model to predict clean audio tokens of DAC codec (Kumar et al. 2024) conditioned on the degraded mel-spectrogram. We use the pre-trained checkpoint in the original work, which comprises $\sim$360M parameters and is trained on a large-scale dataset comprising 20k hours of speech and singing data.
Autoregressive (AR) Model. Our model for AR paradigm, which we term AR+Soundstorm, follows a two-stage pipeline inspired by the architecture introduced in MaskGCT (Wang et al. 2025b; Zhang et al. 2024c). The first stage employs a 0.5B parameter autoregressive model, initialized from the weights of Qwen2.5-0.5B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-0.5B-Instruct. This model takes the noisy audio’s dense features from a pre-trained w2v-bert-2 model [^4] https://huggingface.co/facebook/w2v-bert-2.0 at the 17th-layer as input and is trained to predict the semantic tokens of clean audio extracted from MaskGCT’s semantic codec [^5] https://huggingface.co/amphion/MaskGCT/tree/main/semantic_codec. We use MaskGCT’s second stage [^6] https://huggingface.co/amphion/MaskGCT/tree/main/s2a_model, which is a non-autoregressive Soundstorm-based model (Borsos et al. 2023) with approximately 300M parameters. This model converts the predicted semantic tokens into acoustic tokens, which are then synthesized into the final clean audio waveform.
Flow-Matching (FM) Model. Our FM model, named Flow-SR, is designed to learn a continuous mapping from degraded to clean audio representations. It follows the Diffusion Transformer (DiT) (Peebles and Xie 2023) architecture, which adapts transformer-based models for continuous-time generative tasks. The core of the model consists of 10 LLaMA-2-style decoder layers with 1024 hidden dimensions and 16 attention heads, comprising approximately 220M parameters. The model is also conditioned on dense features of noisy audios from the 17th layer of the pre-trained w2v-bert-2 model and is trained using a flow-matching objective (Lipman et al. 2022) to predict the clean mel-spectrogram (at a 24kHz sampling rate). For the final waveform synthesis, we employ a pre-trained Vocos vocoder (Siuzdak 2023), specifically leveraging the open-source Vevo1.5 implementation [^7] https://huggingface.co/amphion/Vevo1.5/tree/main/acoustic_modeling/Vocoder (Zhang et al. 2025c; Li et al. 2025c).
B.2 Training Details
Pre-Training Data: For AR+Soundstorm and Flow-SR, we sampled 7k hours of clean speech from the Emilia dataset (He et al. 2024). And we use the same noise & rir introduced in AnyEnhance (Zhang et al. 2025a) which contains $\sim$1.2k hours noise and 62,668 rir files, to generate the noisy and reverberant audios.
Pre-Training Procedure: We both train AR+Soundstorm and Flow-SR using the Adam (Kingma and Ba 2014) optimizer, with a learning rate of 5e-4 and 1e-4 respectively and 32K warmup steps, following the inverse square root learning schedule. The training is performed on 4 GPUs, with each GPU processing 5000 frames of audio per step, which is equivalent to 100 seconds of audio, for 200k steps. The training takes approximately 2.5 days for both two models.
DPO-Training Procedure: We also use Adam optimizer for DPO training, with different hyperparameters for different paradigms, all models are trained on 2 GPUs:
  • AnyEnhance: We use a learning rate of 1e-6 with a warmup of 4k steps and a batch size of 4 pairs of samples (i.e., one positive and one negative sample) per GPU, and train for 60k steps. The $\beta$ in DPO loss is set to 10.
  • AR+Soundstorm: We use a learning rate of 1e-5 with a warmup of 4k steps and a batch size of 4 pairs of samples per GPU, and train for 40k steps. The $\beta$ in DPO loss is set to 0.1.
  • Flow-SR: We use a learning rate of 1e-7 with a warmup of 4k steps and a batch size of 12 pairs of samples per GPU, and train for 100k steps. The $\beta$ in DPO loss is set to 1000.
Appendix C Evaluation Details
C.1 Baseline Models
Besides comparison with the three models before and after DPO alignment, we also compare with the following baselines:
  • Voicefixer (Liu et al. 2022): A discriminative general speech restoration model, consists of a ResUNet-based analysis stage which predicts a ratio mask over the degraded mel-spectrogram, and a synthesis stage which converts the masked mel-spectrogram into the final waveform using TFGAN-based vocoder. We use its pre-trained checkpoint [^8] https://github.com/haoheliu/voicefixer.
  • MaskSR (Li, Wang, and Liu 2024): A masked generative model for speech restoration tasks, which predicts acoustic tokens from DAC codec (Kumar et al. 2024) conditioned on the noisy mel-spectrogram. We report the results in the original paper.
  • GenSE (Yao et al. 2025a): A two-stage autoregressive generative framework for speech enhancement, which decouples denoising and generation by first mapping noisy speech to clean semantic tokens and then generating clean acoustic tokens. We report the results in the original paper.
  • LLaSE-G1 (Kang et al. 2025): A LLaMA-based generative speech enhancement model that preserves acoustic consistency by predicting fine-grained speech tokens from continuous WavLM features, and supports diverse SE tasks like denoise, echo cancellation, etc. We report the results in the original paper.
  • FlowSE (Wang et al. 2025d): A flow-matching-based speech enhancement model that learns a continuous mapping from noisy mel-spectrogram to clean mel-spectrogram, with optional text input conditioning capability. We report the results in the original paper.
C.2 Evaluation Details
Metrics: We adopt the following objective and subjective metrics, which are designed to assess various aspects of enhancement quality, including perceptual quality, signal fidelity, and content consistency and timbre preservation:
  • DNSMOS (Reddy, Gopal, and Cutler 2022): a reference-free metric for 16 kHz audio, providing SIG, BAK, and OVRL scores (1-5) for signal quality, background noise, and overall quality.
  • NISQA (Mittag et al. 2021): a reference-free metric for 48 kHz audio that outputs a single overall perceptual quality score (1-5).
  • SpeechBERTScore (Saeki et al. 2024): evaluates semantic similarity between enhanced and reference signals using HuBERT-base [^9] https://huggingface.co/facebook/hubert-base-ls960.
  • Similarity: measures speaker similarity via cosine distance between WavLM-based embeddings [^10] https://huggingface.co/microsoft/wavlm-base-plus-sv.
Figure 5: AB Test interface for subjective evaluation.
Subjective Evaluation We conduct subjective evaluation using an AB test, where participants are asked to choose the preferred audio sample between two options. The interface is shown in Figure 5. Each audio pair was evaluated by 10 raters, the majority of whom are researchers working in speech generation and restoration.
Evaluation Testsets. We assess our models on a diverse range of benchmarks covering two primary restoration scenarios: General Speech Restoration (GSR) and Speech Enhancement (SE).
  • General Speech Restoration (GSR): To evaluate performance on complex, simultaneous distortions (e.g., noise, reverberation, clipping, and bandwidth limitation), we use three test sets: the official Voicefixer-GSR dataset (Liu et al. 2022), and two simulated sets introduced in AnyEnhance based on clean speech from LibriVox (Kearns 2014) and clean singing from CCMusic (Zhou et al. 2024).
  • Speech Enhancement (SE): To evaluate performance specifically on downstream SE tasks, we use the official test sets from the DNS Challenge 2020 (Reddy et al. 2020), including both the with-reverb and no-reverb conditions.
Appendix D More Experimental Results
D.1 Full Results on DNS Challenge Benchmarks
Table 8 presents the detailed results on the DNS Challenge benchmarks. Our findings confirm that the DPO-aligned models consistently outperform their original counterparts across both non-reverberant and reverberant conditions. Specifically, we observe significant gains in DNSMOS scores (SIG, BAK, OVRL) and substantial boosts in perceptual quality (NISQA) for all three paradigms. Interestingly, while the alignment strongly favors perceptual and signal quality, there are occasional minor trade-offs in SpeechBERTScore and Speaker Similarity, suggesting the model learns to prioritize clarity and noise reduction, which aligns well with the goals of speech enhancement.
D.2 AB Test on CCMusic-GSR Testset
Figure 6: Human AB preference test results for DPO-aligned models on CCMusic-GSR testset.
To further validate the effectiveness of our DPO-aligned models, we also conducted an AB preference test on the CCMusic-GSR testset. The results, shown in Figure 6, show consistent improvements across all paradigms, with win rates of 30.3% (MGM), 48.5% (AR), and 33.3% (FM), demonstrating the effectiveness of DPO alignment in the singing voice domain.
Figure 7: Spectrogram comparison for the AR+Soundstorm (AR) model. Each row compares the original noisy audio (left), the restored audio from our DPO-aligned model (center), and the output from the baseline model without alignment (right).
D.3 More Result on Multi-Metric Preference Ablation
To fully validate the robustness of our multi-metric alignment strategy, we conducted additional ablation studies on the preference metric for both the FM and MGM models, supplementing the AR model results presented in the main paper. The results, shown in Table 9 and Table 10, demonstrate a consistent pattern across all generative paradigms.
Criterion SIG BAK OVRL NISQA SBERT SIM
- 3.398 3.969 3.104 4.010 0.812 0.918
Multi-Metric 3.483 4.092 3.230 4.672 0.830 0.924
NISQA 3.415 4.035 3.147 4.495 0.814 0.920
OVRL 3.464 4.010 3.179 4.206 0.815 0.918
SBERT 3.413 4.002 3.129 4.209 0.826 0.919
SIM 3.394 3.951 3.091 4.095 0.810 0.921
Table 9: Ablation on preference metric for the FM model on Voicefixer-GSR testset.
Criterion SIG BAK OVRL NISQA SBERT SIM
- 3.406 4.073 3.136 4.308 0.829 0.924
Multi-Metric 3.532 4.091 3.267 4.639 0.834 0.935
NISQA 3.429 4.096 3.177 4.715 0.811 0.930
OVRL 3.536 4.088 3.274 4.459 0.820 0.928
SBERT 3.407 4.031 3.121 4.364 0.861 0.929
SIM 3.372 4.060 3.107 4.381 0.841 0.936
Table 10: Ablation on preference metric for the MGM model on Voicefixer-GSR testset.
As with the AR model, aligning with a single metric (e.g., NISQA or OVRL) on the FM and MGM models improves the targeted metric but often leads to stagnation or even degradation in other, non-targeted quality dimensions. For example, in Table 10, while the SIM-aligned MGM model improves speaker similarity (SIM), its signal quality scores (SIG and OVRL) fall below the original baseline. In contrast, the Multi-Metric alignment strategy consistently yields balanced and comprehensive improvements across nearly all evaluation metrics for both models. This reinforces our core hypothesis that a holistic preference signal, derived from a unanimous agreement across a diverse suite of metrics, is crucial for mitigating reward hacking and achieving genuinely superior restoration quality, regardless of the underlying generative architecture.
D.4 Application: Empowering Discriminative Models via Pseudo-Labeling
To showcase the practical utility of our aligned models, we explore their potential as “data annotators” in a challenging, real-world scenario: singing voice restoration, where paired clean/noisy data is often scarce.
Dataset. We use a proprietary dataset of 183 song recordings. The dataset features 10 different singers and was captured with a variety of recording devices (including smartphones and condenser microphones) across diverse acoustic environments such as piano rooms, rehearsal halls, concert halls, classrooms, bedrooms, and even reverberant spaces like hallways, stairwells, and basements. We partitioned this dataset into a training set of 170 songs (6.97 hours, 3,121 audio clips) and a test set of 13 songs (0.80 hours, 416 audio clips).
Figure 8: Spectrogram comparison for the AnyEnhance (MGM) model. Each row compares the original noisy audio (left), the restored audio from our DPO-aligned model (center), and the output from the baseline model without alignment (right).
Pseudo-Label Generation and Fine-tuning. We first prepared a specialized AnyEnhance model, which was pre-trained with additional singing voice data via the SingNet pipeline (Gu et al. 2025) and then aligned on singing voice data with our multi-metric preference alignment strategy. This expert generative model was then used to process the noisy recordings from our training set, generating high-quality pseudo-clean labels for each segment.
Subsequently, we fine-tuned the official Voicefixer checkpoint on these pseudo-labeled pairs. The fine-tuning was conducted for 5k steps using the Adam optimizer with a learning rate of 1e-5 and a batch size of 24. As shown in Section 4.4 in the main paper, this approach significantly boosts the performance of the original Voicefixer model. Visualizations of the enhancement quality can be seen in Figure 10.
D.5 Audio Samples Visualization
To provide a qualitative and visual demonstration of our alignment strategy’s impact, we present spectrograms of several audio samples in Figures 7, 8, and 9. Each figure corresponds to one of the three generative paradigms. Within each figure, every row showcases a direct comparison between the original noisy input (left column), the output from our DPO-aligned model (center column), and the output from the baseline model without alignment (right column).
As can be observed across all examples, the aligned models (w/ Align) consistently exhibit superior performance in suppressing background noise, which appears as a ”haze” in the spectrograms of the noisy and baseline outputs. Furthermore, our aligned models produce much clearer and more defined harmonic structures (the bright horizontal lines), indicating a higher fidelity restoration of the speech content. This visual evidence strongly corroborates the quantitative improvements reported in the main paper.
Finally, Figure 10 visualizes the results from our application experiment on singing voice restoration. This figure demonstrates the practical utility of our aligned models as powerful “data annotators” for training discriminative models in data-scarce scenarios, as the fine-tuned Voicefixer model (center column) significantly outperforms the original Voicefixer model (right column) in terms of clarity and artifact removal.
Figure 9: Spectrogram comparison for the Flow-SR (FM) model. Each row compares the original noisy audio (left), the restored audio from our DPO-aligned model (center), and the output from the baseline model without alignment (right).
Figure 10: Visualization of the singing voice restoration application, demonstrating our aligned model as a “data annotator”. Each row compares the raw recording (left), the output from the Voicefixer model fine-tuned on our pseudo-labels (center), and the output from the original, un-tuned Voicefixer model (right). The fine-tuned version shows remarkable improvements in clarity and artifact removal.